MMA tests the hypothesis that the average nbr of FNP is more than 4. Consider this to help carry out the test. n=30, x̅=5.5, σ=3 and a=0.01. Answer Q 12 to 16

Q 12
The correct set of hypothesis is
1. Ho : μ = 4; H1 : μ ≠ 4
2. Ho : μ = 5.5; H1 : μ ≠ 5.5
3. Ho : μ = 4; H1 : μ > 4
4. Ho : μ = 5.5; H1 : μ > 5.5
5. Ho : X = 4; H1 : X > 4

Q 13
The applicable test statistic is
1. Z test
2. t test
3. Chi–Square test
4. F–test

Q 14
The value of the test statistic and the critical value are
1. 1.5 and 2.33
2. 2.74 and 2.33
3. 2.74 and 2.58
4. 1.5 and 2.58

Q 15
The p-value for the test is
1. 0.0062
2. 0.0668
3. 0.1336
4. 0.0031

Q 16
What decision and conclusion can we take with regard to the hypothesis test
1. Reject null hypothesis(NH) and conclude that the average nbr of FNP > 4
2. Reject NH and conclude that the average nbr of FNP = 4
3. Don't eject NH and conclude that the average nbr of FNP = 4
4. Accept the NH
1
Expert's answer
2020-05-11T17:38:26-0400

Q 12

The correct set of hypothesis is 

3. Ho : μ = 4; H1 : μ > 4 


Q 13

The applicable test statistic is

1. Z test 


Q 14

The value of the test statistic and the critical value are 


"z_c=z_{1-\alpha}=2.33"

"z={\bar{x}-\mu\over \sigma/\sqrt{n}}={5.5-4\over 3/\sqrt{30}}\approx2.7386"

2. 2.74 and 2.33 


Q 15

The p-value for the test is "p=0.003085<0.01"

4. 0.0031 


Q 16

What decision and conclusion can we take with regard to the hypothesis test 

Since "p=0.003085<0.01," it is concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is greater than 4, at the 0.01 significance level.

1. Reject null hypothesis(NH) and conclude that the average nbr of FNP > 4